# Massively Parallel Proof-Number Search for Impartial Games and Beyond

### v1

@@ Introduction
Game-solving is a well-established and difficult task in artificial intelligence, which involves computing outcomes under perfect play of all players, often requiring a deep traversal of vast game trees. Despite these challenges, several classical games have already been solved, including Connect Four (allis1988connect4), Gomoku (allis1996gomoku), Checkers (schaeffer2007checkers), and Othello (takizawa2024othello). Proof-Number Search (allis1994) is among the most successful game-solving algorithms. It is a best-first tree search algorithm designed to compute game outcomes efficiently. This algorithm has been widely used due to its ability to focus on the most promising parts of the game tree, making it particularly effective in domains with large branching factors or unbalanced trees. Beyond games, its variants and related algorithms have been applied in various other settings, for example, in chemistry (franz2022), graphical models (dechter2007), medicine (heifets2012), and theorem proving (lample2022).
With increasing computational power, it is natural to pursue parallel versions of Proof-Number Search; yet, despite many efforts, existing variants have failed to scale efficiently across many CPUs. Several studies have suggested using parallel Proof-Number Search or its variants in diverse domains (franz2022; kishimoto2015), but scalability remains a challenge. These limitations led kishimoto2012gametree to pose the problem of developing a well-scaling, massively parallel Proof-Number Search.
Many combinatorial games remain far from being solved, including Chess, Shogi, Go, Cram, or Sprouts. In Sprouts, two players alternately connect $n$ given spots in the plane according to simple rules, where the last player unable to make a move loses; see cizek2021implementation for the implementation. We use this game as our experimental domain, as it features highly unbalanced game trees with a large branching factor, which makes it well-suited for Proof-Number Search. Designed by Conway and Paterson to resist computer analysis (roberts2015genius), Sprouts poses a challenging benchmark: the complexity of its game tree surpasses Chess and Go for relatively small values of $n$; see Figure 1. The computation of outcomes of Sprouts positions is further motivated by the long-standing Sprouts Conjecture (applegate1991computer), which states that the $n$-spot position, consisting of $n$ given spots, is winning for the first player if and only if $n$ is congruent to 3, 4, or 5 modulo 6.
@@
We address the problem posed by kishimoto2012gametree and develop a new variant of Proof-Number Search that scales efficiently on many CPUs. Building on this, we implement a parallel solver for combinatorial games. Together with a new memory-efficient algorithm for impartial games, we achieve numerous new results for the game Sprouts.
@@
Proof-Number Search is designed to solve problems that are represented as AND/OR trees. We formalize the notion of game trees with Grundy numbers to reduce the complexity of game trees of the so-called impartial games, as introduced by lemoineViennot12Inevitable. We adapt Depth-First Proof-Number Search with Grundy numbers, a popular and memory-efficient variant of Proof-Number Search, to work with such trees.
@@
To demonstrate the efficiency of our new algorithms, we implemented a solver for the Sprouts game, and used it to determine many new outcomes. Our solver outperforms the state-of-the-art solver GLOP by lemoine2015nimber, achieving a four-order-of-magnitude speedup and generating proofs 1,000 times more complex. Despite the exponential growth of the game trees, our solver verifies the Sprouts Conjecture for 42 new positions, nearly doubling the number of known outcomes.
@@
A considerable amount of work has been devoted to designing parallel versions of Proof-Number Search. These include RP-PNS, introduced by saito2010parallel, a parallel variant of DFPN by kaneko2010parallel, and SPDFPN algorithm by pawlewicz2013scalable. All of these were designed for shared-memory systems, which typically offer only a limited number of CPU cores. To leverage the greater computational power of clusters with many cores, algorithms tailored to distributed-memory systems are required. Such algorithms include ParaPDS by kishimoto1999parallel, Job-Level Proof-Number Search by wu2010joblevel; wu2013joblevel, and PPN2 search by saffidine2012parallel. A distributed-memory parallel Proof-Number Search was also used by schaeffer2005solving; schaeffer2007checkers in solving Checkers.
As noted by kishimoto2012gametree, the scalability of existing algorithms was evaluated only on a relatively small number of cores, and, despite various successes in game solving, their efficiency declined rapidly as core counts increased. While parallel DFPN by kaneko2010parallel and SPDFPN have speedups of 3.58 on 8 cores and 11.8 on 16 cores, respectively, they are limited to shared-memory systems. The best scaling factors for distributed-memory systems are reported for PPN2 search, which achieved speedups of 7.2, 11.1, and 18.3 on 16, 32, and 64 cores, respectively.
Determining the outcomes of $n$-spot Sprouts positions has a long history dating back to the 1960s. Conway himself determined the outcomes for $n\leq 3$, and later Mollison extended the results to $n=6$ by hand (gardner1967mathematical). The first computer-based solver was developed by applegate1991computer, who used it to prove outcomes for $n\leq 11$. In 2006, Purinton created the program AuntBeast, which solved positions up to 14 spots. A major advance was made by the state-of-the-art solver GLOP (lemoine2015nimber) who used Alpha-Beta Pruning combined with Grundy number theory to solve all positions for $n\leq 32$. These results were later extended to all values of $n\leq 44$ and selected cases up to $n=53$, by incorporating a basic variant of Proof-Number Search combined with Grundy numbers. All known outcomes agree with the Sprouts Conjecture.
@@
We now briefly review relevant preliminaries from combinatorial game theory; for more details, see kishimoto2012gametree and berlekamp2003winning. A combinatorial game is a two-player, deterministic game with perfect information that ends in a finite number of moves. Under normal play convention, the last player able to move wins. A combinatorial game is impartial if the available moves from any given position are the same for both players, regardless of whose turn it is. Combinatorial games include Chess, Checkers, Go, Othello, and Shogi, whereas impartial games include Nim, Sprouts, Quarto, and Cram.
@@
Game trees of combinatorial games can be represented by the so-called AND/OR trees. Here, we instead work with AND/OR trees in negamax fashion using NAND trees, which are easier to work with, as they eliminate the unnecessary distinction between the AND and OR nodes in subsequent definitions. The nodes at odd and even levels of a NAND tree correspond to positions where the first and second player, respectively, is on the move. In particular, the root corresponds to the initial position with the first player on the move. A terminal is a node with no children that represents a position where the game ends. In a partially expanded NAND tree, a node with at least one child is internal if its children are generated; otherwise, it is a leaf.
@@
Given a position $P$, the goal is to find a proof of $P$, which is a NAND tree rooted in $P$ whose value is known. These proofs correspond to so-called weak solutions, as they give the outcome of $P$ and a winning strategy for one of the players. In contrast, strong solutions provide outcomes and winning strategies for all positions reachable in the game.
@@
We first describe a basic variant PNS of the Proof-Number Search for NAND trees. In each step, PNS selects and expands a most proving node (MPN)—a leaf in a currently expanded tree whose solution would contribute the most to the solution of the root. To find an MPN, each node $v$ in a NAND tree is associated with a proof number $pn(v)\in\mathbb{N}_{0}\cup\{\infty\}$ and a disproof number $dn(v)\in\mathbb{N}_{0}\cup\{\infty\}$, representing the lower bounds on the minimum number of leaves in the subtree of $v$ that have to be solved to prove or disprove $v$, respectively. Both these values are computed in a bottom-up manner. For leaves, they are initialized as $pn(v)=dn(v)=1$. For terminals in games under normal play, $pn(v)=\infty$ and $dn(v)=0$. The values $pn(v)$ and $dn(v)$ of an internal node $v$ are defined as $pn(v)=\min_{c}{dn(c)}$ and $dn(v)=\sum_{c}{pn(c)}$, where the minimum and the sum are taken over the children $c$ of $v$. At the start, the algorithm initializes $pn(r)$ and $dn(r)$ of the root $r$. At each step, it then finds an MPN by descending from the root and always selecting the child with the lowest disproof number with ties broken arbitrarily. PNS then expands the selected MPN by generating all its children $c$ and initializing their $pn(c)$ and $dn(c)$. At the end of each step, it updates the proof and disproof numbers on the path back to the root.
@@
One drawback of PNS is that it stores the entire expanded tree in memory, which is usually consumed very quickly. Therefore, nagai1999phdthesis introduced Depth-first Proof-Number Search (DFPN), a space-efficient variant of PNS.
@@
We initialize $pt(r)=dt(r)=\infty$ for the root $r$. Let $v$ be the currently visited node, $w$ and $w^{\prime}$ be its children with the lowest and second lowest disproof number, where $w$ is the next node to be selected. The thresholds of $w$ are then set as
@@
As shown by lemoineViennot12Inevitable and by belingRogalski2020, the Sprague–Grundy Theorem (grundy1939mathematics; sprague1935uber) can be applied to effectively simplify game trees of impartial games under the normal play convention; see Figure 1. Since we build on these techniques, we briefly recall the necessary background.
Nim is a game that is played on $h$ heaps with $n_{1},\dots,n_{h}\in\mathbb{N}_{0}$ objects. With each move, a player must remove at least one object from a single heap, and the first player with no move loses the game. Nim is a strongly solved game, as bouton1901nim proved that the outcome of Nim is a loss if and only if $n_{1}\oplus\cdots\oplus n_{h}=0$, where $\oplus$ is a bitwise exclusive or.
Let $P_{1},\dots,P_{k}$ be positions of an impartial game under the normal play convention. Then, the combination of $P_{1},\dots,P_{k}$, denoted by $P_{1}+\dots+P_{k}$, is a position in which, in each turn, the players decide to move in one of the positions $P_{1},\dots,P_{k}$ while leaving the other positions untouched. The first player with no move in any of $P_{1},\dots,P_{k}$ loses in the combination $P_{1}+\dots+P_{k}$. A position $P$ is atomic if it cannot be expressed as a combination of at least two non-empty positions, otherwise it is decomposable, in which case $P=P_{1}+\dots+P_{k}$ for non-empty atomic positions $P_{1},\dots,P_{k}$ and $k\geq 2$. Two positions $P$ and $Q$ are equivalent if, for any position $R$, the combinations $P+R$ and $Q+R$ have the same outcome.
The Grundy number $gn(P)$ of a position $P$ (also called the nimber of $P$) is defined recursively as follows. If $P$ is a terminal position, then $gn(P)=0$. Otherwise, $gn(P)$ is equal to $\min{\mathbb{N}_{0}\setminus G}$ where $G$ is the set of the Grundy numbers of the children of $P$. Using $\ast n$ to denote a Nim position with a single heap of $n$ objects, we formulate the Sprague–Grundy Theorem (grundy1939mathematics; sprague1935uber), which states that each equivalence class of impartial games under the normal play convention can be represented by a unique Grundy number.
@@
It follows from Theorem 1 that the outcome of $P_{1}+\dots+P_{k}$ is a loss if and only if $gn(P_{1})\oplus\dots\oplus gn(P_{k})=0$. Thus, we can compute the outcome of a decomposable position $P=P_{1}+\cdots+P_{k}$ more efficiently using the Grundy numbers $gn(P_{1}),\dots,gn(P_{k})$ (lemoineViennot12Inevitable).
@@ Methods
First, we formalize extended NAND trees with Grundy numbers that were implicitly used by lemoineViennot12Inevitable to efficiently determine the outcomes of decomposable positions. Then, we describe and improve PNS for extended NAND trees by lemoine2011phdthesis and introduce DFPN for such trees. Finally, we describe our massively parallel version PNS-PDFPN of the Proof Number Search.
@@
To apply the Sprague–Grundy Theorem, we need to determine the Grundy number $gn(P)$ of a position $P$ in an impartial game $G$ under the normal play convention. We efficiently compute $gn(P)$ using the following recursive procedure introduced by lemoineViennot12Inevitable: start with $n=0$ and increment $n$ until the outcome of the combination $P+\ast n$, called a couple, is loss. Then $gn(P)=n$, because $gn(P)=n$ if and only if the outcome of $P+\ast n$ is loss.
@@
We define the proof and disproof numbers so that they give lower bounds on the size of the proof of a given position, which guarantees that PNS finds a minimal proof. For leaves, we initialize them to $1$. For atomic nodes, the proof and disproof numbers and the next node $w$ are defined as in the original setting. If $u$ is a Grundy node, then it corresponds to an atomic position $Q$ and we use $v_{i}$ to denote the child of $u$ representing $Q+\ast i$. The children of $u$ are atomic and are generated incrementally with increasing $i$, each only after the previous one has been proved, as prescribed by the recursive procedure for the calculation of $gn(Q)$. If $v_{j}$ is the last child of $u$ generated so far, then we set $pn(u)=dn(u)=\min{\{pn(v_{j}),dn(v_{j})\}}$ and $w=v_{j}$. Finally, let $v$ be a decomposable node. Then $v$ corresponds to a couple $P+\ast n$ where $P=P_{1}+\cdots+P_{k}$ and $k\geq 2$. We use $u_{1},\dots,u_{k}$ to denote the nodes where each $u_{i}$ is a Grundy node corresponding to $P_{i}$ and $v^{\prime}_{k}$ is an atomic node corresponding to $P_{k}+\ast n^{\prime}$. Note that only $u_{1},\dots,u_{k-1},v^{\prime}_{k}$ are the children of $v$, and that in order to generate $v^{\prime}_{k}$, we first need to know the Grundy numbers of $P_{1},\dots,P_{k-1}$, as $n^{\prime}=n\oplus gn(P_{1})\oplus\cdots\oplus gn(P_{k-1})$. Thus, if $v^{\prime}_{k}$ has not yet been generated, we set $pn(v)=dn(v)=\sum_{i=1}^{k-1}pn(u_{i})$ and $w$ to be the first node from $u_{1},\dots,u_{k-1}$ whose Grundy number is unknown. Otherwise, we let $pn(v)=pn(v^{\prime}_{k})$, $dn(v)=dn(v^{\prime}_{k})$, and $w=v^{\prime}_{k}$.
We improve this version of PNS from lemoine2011phdthesis by following schijf1994transpositions and merging nodes representing identical states to transform the tree into a directed acyclic graph. This significantly reduces the computation time without increasing memory overhead when combined with storing all touched nodes.
@@
To simplify the definition of the thresholds, we replace each Grundy node $u$ in its decomposable parent with the node $v_{j}$, which would be selected as the next node if we were in $u$. Then, we can keep the thresholds only for atomic and decomposable nodes. Now, let $v$ be the currently visited node, and let $w$ be the next node to be selected. For atomic $v$, we let $w^{\prime}$ be its child with the second-lowest disproof number and we set the thresholds as
@@
where ps(w)=ds(v)+dn(v)pn(w)ps(w)=ds(v)+dn(v)-pn(w) and $ds(w)=ps(v)$. For decomposable $v$, we set $pt(w)=pt(v)$, $dt(w)=dt(v)$, and $mt(w)=mt(v)$ if $v^{\prime}_{k}$ has been generated. If $v^{\prime}_{k}$ has not been generated yet, we define the thresholds as
@@
where $ps(w)=ds(w)=0$ and $t(v)=\min\{pt(v),dt(v),\allowbreak mt(v)-\min\{ps(v),ds(v)\}\}$. For the root $r$, we initialize $pt(r)=dt(r)=mt(r)=\infty$ and $ps(r)=ds(r)=0$. With this choice of the thresholds, we obtain the following result.
@@
See Appendix for the proof. To balance the time complexity and memory consumption of DFPN with GN, we also incorporate a transposition table of proof and disproof numbers, maintained by the replacement strategy proposed by nagai1999phdthesis. In addition, we store the Grundy numbers derived during computation in a separate database. Since this database is much smaller and fits easily into memory, we do not apply any replacement strategy to it.
@@
Our new parallel variant PNS-PDFPN of Proof-Number Search operates on two parallelized levels. The first PNS level targets distributed-memory systems and is based on Job-Level Proof-Number Search used by schaeffer2007checkers, saffidine2012parallel, and wu2013joblevel. At this level, the master process maintains the current proof state and repeatedly assigns jobs to workers for asynchronous second-level processing. At the second level, each worker performs the parallel DFPN algorithm by kaneko2010parallel, which we call PDFPN, to utilize shared memory within a single cluster node. We introduce the sharing of key results between workers to reduce search overhead. We now describe both levels in detail.
@@
The master and workers typically run on separate nodes of a cluster and communicate over an interconnected network. When a worker becomes idle, the master selects a so-called pseudo-MPN leaf $\ell$ and assigns it to the worker. The worker then processes $\ell$ asynchronously until it is solved or the maximum number of expansions is reached. The resulting proof and disproof numbers of $\ell$ and its children are then sent back to the master, which uses them to expand $\ell$ in the master tree and updates its proof and disproof numbers. During processing of $\ell$, the worker periodically sends updated values $pn(\ell)$ and $dn(\ell)$ to keep the master tree as up-to-date as possible. To prevent reassignment of the same jobs, the assigned leaves are locked, and the definition of proof and disproof numbers is slightly adjusted to avoid misleading attraction to the locked leaves; see wu2013joblevel for details.
Similar to xichen15Hex and their JL-UCT Search, our master maintains a database of key computed results shared with all workers to reduce search overhead. Newly derived results are sent to workers along with their assigned jobs, while workers report new results back to the master with each update. Finally, if multiple workers are located on the same node, then they are grouped together to share the same local database.
@@
Our workers are based on the DFPN algorithm to make full use of available memory; thus, we parallelize using PDFPN. In this algorithm, there are as many threads as the number of cores assigned to the worker. Each thread performs an independent DFPN search in the subtree of $\ell$, while sharing the lock-protected transposition table of proof and disproof numbers. To decrease redundant thread computation in shared subtrees, the disproof numbers are virtually increased during the selection of the next node $w$ by the number of threads currently computing in the corresponding subtrees. The thresholds in (1) and (2) are then modified by subtracting $th(w)$ from the term $dn(w^{\prime})+1$, where $th(w)$ is the number of threads in the subtree of $w$. Additionally, if a thread solves a position currently computed by a different one, the other thread is notified to backtrack to that position.

### v2

@@ Introduction
Game-solving is a well-established and difficult task in artificial intelligence, which involves computing outcomes under perfect play of all players, often requiring a deep traversal of vast game trees. Despite these challenges, several classical games have already been solved, including Connect Four (Allis 1988), Gomoku (Allis et al. 1996), Checkers (Schaeffer et al. 2007), and Othello (Takizawa 2023). Proof-Number Search (Allis et al. 1994) is among the most successful game-solving algorithms. It is a best-first tree search algorithm designed to compute game outcomes efficiently. This algorithm has been widely used due to its ability to focus on the most promising parts of the game tree, making it particularly effective in domains with large branching factors or unbalanced trees. Beyond games, its variants and related algorithms have been applied in various other settings, for example, in chemistry (Franz et al. 2022), graphical models (Dechter and Mateescu 2007), medicine (Heifets and Jurisica 2012), and theorem proving (Lample et al. 2022).
With increasing computational power, it is natural to pursue parallel versions of Proof-Number Search; yet, despite many efforts, existing variants have failed to scale efficiently across many CPUs. Several studies have suggested using parallel Proof-Number Search or its variants in diverse domains (Franz et al. 2022; Kishimoto et al. 2015), but scalability remains a challenge. These limitations led Kishimoto et al. (2012) to pose the problem of developing a well-scaling, massively parallel Proof-Number Search.
Many combinatorial games remain far from being solved, including Chess, Shogi, Go, Cram, or Sprouts. In Sprouts, two players alternately connect $n$ given spots in the plane according to simple rules, where the last player unable to make a move loses; see Čížek and Balko (2021) for the implementation. We use this game as our experimental domain, as it features highly unbalanced game trees with large branching factors, making it well-suited for Proof-Number Search. Designed by Conway and Paterson to resist computer analysis (Roberts 2015), Sprouts poses a challenging benchmark: the complexity of its game tree surpasses Chess and Go for relatively small values of $n$; see Figure 1. The computation of outcomes of Sprouts positions is further motivated by the long-standing Sprouts Conjecture (Applegate et al. 1991), which states that the $n$-spot position, consisting of $n$ given spots, is winning for the first player if and only if $n$ is congruent to 3, 4, or 5 modulo 6.
@@
We address the problem posed by Kishimoto et al. (2012) and develop a new variant of Proof-Number Search that scales efficiently on many CPUs. Building on this, we implement a parallel solver for combinatorial games. Together with a new memory-efficient algorithm for impartial games, we achieve numerous new results for the game Sprouts.
@@
Proof-Number Search is designed to solve problems that are represented as AND/OR trees. We formalize the notion of game trees with Grundy numbers to reduce the complexity of game trees of the so-called impartial games, as introduced by Lemoine and Viennot (2012). We adapt Depth-First Proof-Number Search with Grundy numbers, a popular and memory-efficient variant of Proof-Number Search, to work with such trees.
@@
To demonstrate the efficiency of our new algorithms, we implemented a solver for the Sprouts game, and used it to determine many new outcomes. Our solver outperforms the state-of-the-art solver GLOP by Lemoine and Viennot (2015), achieving a four-order-of-magnitude speedup and generating proofs 1,000 times more complex. Despite the exponential growth of the game trees, our solver verifies the Sprouts Conjecture for 42 new positions, nearly doubling the number of known outcomes.
@@
A considerable amount of work has been devoted to designing parallel versions of Proof-Number Search. These include RP-PNS, introduced by Saito et al. (2010), a parallel variant of DFPN by Kaneko (2010), and SPDFPN algorithm by Pawlewicz and Hayward (2014). All of these were designed for shared-memory systems, which typically offer only a limited number of CPU cores. To leverage the greater computational power of clusters with many cores, algorithms tailored to distributed-memory systems are required. Such algorithms include ParaPDS by Kishimoto and Kotani (1999), Job-Level Proof-Number Search by Wu et al. (2011, 2013), and PPN2 search by Saffidine et al. (2012). A distributed-memory parallel Proof-Number Search was also used by Schaeffer et al. (2005, 2007) in solving Checkers.
As noted by Kishimoto et al. (2012), the scalability of existing algorithms was evaluated only on a relatively small number of cores, and, despite various successes in game solving, their efficiency declined rapidly as core counts increased. While parallel DFPN by Kaneko (2010) and SPDFPN have speedups of 3.58 on 8 cores and 11.8 on 16 cores, respectively, they are limited to shared-memory systems. The best scaling factors for distributed-memory systems are reported for PPN2 search, which achieved speedups of 7.2, 11.1, and 18.3 on 16, 32, and 64 cores, respectively.
Determining the outcomes of $n$-spot Sprouts positions has a long history dating back to the 1960s. Conway himself determined the outcomes for $n\leq 3$, and later Mollison extended the results to $n=6$ by hand (Gardner 1967). The first computer-based solver was developed by Applegate et al. (1991), who used it to prove outcomes for $n\leq 11$. In 2006, Purinton created the program AuntBeast, which solved positions up to 14 spots. A major advance was made by the state-of-the-art solver GLOP (Lemoine and Viennot 2015) who used Alpha-Beta Pruning combined with Grundy number theory to solve all positions for $n\leq 32$. These results were later extended to all values of $n\leq 44$ and selected cases up to $n=53$, by incorporating a basic variant of Proof-Number Search combined with Grundy numbers. All known outcomes agree with the Sprouts Conjecture.
@@
We now briefly review relevant preliminaries from combinatorial game theory; for more details, see Kishimoto et al. (2012) and Berlekamp et al. (2003). A combinatorial game is a two-player, deterministic game with perfect information that ends in a finite number of moves. Under normal play convention, the last player able to move wins. A combinatorial game is impartial if the available moves from any given position are the same for both players, regardless of whose turn it is. Combinatorial games include Chess, Checkers, Go, Othello, and Shogi, whereas impartial games include Nim, Sprouts, Quarto, and Cram.
@@
Game trees of combinatorial games can be represented by the so-called AND/OR trees. Here, we instead work with AND/OR trees in negamax fashion using NAND trees, which are easier to work with, as they eliminate the unnecessary distinction between the AND and OR nodes in subsequent definitions. The nodes at odd and even levels of a NAND tree correspond to positions where the first and second player, respectively, is on the move. In particular, the root corresponds to the initial position with the first player on the move. A terminal is a node with no children, which represents a position where the game ends. In a partially expanded NAND tree, a non-terminal node is internal if some of its children are generated, and a leaf if none of them are.
@@
Given a position $P$, the goal is to find a proof of $P$, which is a NAND tree rooted in $P$ whose value is known. These proofs correspond to the so-called weak solutions, as they give the outcome of $P$ and a winning strategy for one of the players. In contrast, strong solutions provide outcomes and winning strategies for all positions reachable in the game.
@@
We first describe a basic variant PNS of the Proof-Number Search for NAND trees. In each step, PNS selects and expands a most proving node (MPN)—a leaf in a currently expanded tree whose solution would contribute the most to the solution of the root. To find an MPN, each node $v$ in a NAND tree is associated with a proof number $pn(v)\in\mathbb{N}_{0}\cup\{\infty\}$ and a disproof number $dn(v)\in\mathbb{N}_{0}\cup\{\infty\}$, representing the lower bounds on the minimum number of leaves in the subtree of $v$ that have to be solved to prove or disprove $v$, respectively. Both these values are computed in a bottom-up manner. For leaves, they are initialized as $pn(v)=dn(v)=1$. For terminals in games under the normal play convention, $pn(v)=\infty$ and $dn(v)=0$. The values $pn(v)$ and $dn(v)$ of an internal node $v$ are defined as $pn(v)=\min_{c}{dn(c)}$ and $dn(v)=\sum_{c}{pn(c)}$, where the minimum and the sum are taken over the children $c$ of $v$. At the start, the algorithm initializes $pn(r)$ and $dn(r)$ of the root $r$. At each step, it then finds an MPN by descending from the root and always selecting the child with the lowest disproof number with ties broken arbitrarily. PNS then expands the selected MPN by generating all its children $c$ and initializing their $pn(c)$ and $dn(c)$. At the end of each step, it updates the proof and disproof numbers on the path back to the root.
@@
One drawback of PNS is that it stores the entire expanded tree in memory, which is usually consumed very quickly. Therefore, Nagai (2002) introduced Depth-first Proof-Number Search (DFPN), a space-efficient variant of PNS.
@@
First, set $pt(r)=dt(r)=\infty$ for the root $r$. Let $v$ be the currently visited node, $w$ and $w^{\prime}$ be its children with the lowest and second lowest disproof number, where $w$ is the next node to be selected. The thresholds of $w$ are then set as
@@
As shown by Lemoine and Viennot (2012) and by Beling and Rogalski (2020), the Sprague–Grundy Theorem (Grundy 1939; Sprague 1935) can be applied to effectively simplify game trees of impartial games under the normal play convention; see Figure 1. Since we build on these techniques, we briefly recall the necessary background.
Nim is a game that is played on $h$ heaps with $n_{1},\dots,n_{h}\in\mathbb{N}_{0}$ objects. With each move, a player must remove at least one object from a single heap, and the first player with no move loses the game. Nim is a strongly solved game, as Bouton (1901) proved that the outcome of Nim is loss if and only if $n_{1}\oplus\cdots\oplus n_{h}=0$, where $\oplus$ is a bitwise exclusive or.
Let $P_{1},\dots,P_{k}$ be positions of an impartial game under the normal play convention. Then, the combination of $P_{1},\dots,P_{k}$, denoted by $P_{1}+\dots+P_{k}$, is a position in which, in each turn, the players decide to move in one of the positions $P_{1},\dots,P_{k}$ while leaving the other positions untouched. The first player with no move in any of $P_{1},\dots,P_{k}$ loses in the combination $P_{1}+\dots+P_{k}$. A position $P$ is atomic if it cannot be expressed as a combination of at least two non-empty positions; otherwise, it is decomposable, in which case $P=P_{1}+\dots+P_{k}$ for non-empty atomic positions $P_{1},\dots,P_{k}$ and $k\geq 2$. Two positions $P$ and $Q$ are equivalent if, for any position $R$, the combinations $P+R$ and $Q+R$ have the same outcome.
The Grundy number $gn(P)$ of a position $P$ (also called the nimber of $P$) is defined recursively as follows. If $P$ is a terminal position, then $gn(P)=0$. Otherwise, $gn(P)$ is equal to $\min{\mathbb{N}_{0}\setminus G}$ where $G$ is the set of the Grundy numbers of the children of $P$. Using $\ast n$ to denote a Nim position with a single heap of $n$ objects, we formulate the Sprague–Grundy Theorem (Grundy 1939; Sprague 1935), which states that each equivalence class of impartial games under the normal play convention can be represented by a unique Grundy number.
@@
It follows from Theorem 1 that the outcome of $P_{1}+\dots+P_{k}$ is loss if and only if $gn(P_{1})\oplus\dots\oplus gn(P_{k})=0$. Thus, we can compute the outcome of a decomposable position $P=P_{1}+\cdots+P_{k}$ more efficiently using the Grundy numbers $gn(P_{1}),\dots,gn(P_{k})$ (Lemoine and Viennot 2012).
@@ Methods
First, we formalize extended NAND trees with Grundy numbers that were implicitly used by Lemoine and Viennot (2012) to efficiently determine the outcomes of decomposable positions. Then, we describe and improve PNS for extended NAND trees by Lemoine (2011) and introduce DFPN for such trees. Finally, we describe our massively parallel version PNS-PDFPN of Proof Number Search.
@@
To apply the Sprague–Grundy Theorem, we need to determine the Grundy number $gn(P)$ of a position $P$ in an impartial game $G$ under the normal play convention. We efficiently compute $gn(P)$ using the following recursive procedure introduced by Lemoine and Viennot (2012): start with $n=0$ and increment $n$ until the outcome of the combination $P+\ast n$, called a couple, is loss. Then $gn(P)=n$, because $gn(P)=n$ if and only if the outcome of $P+\ast n$ is loss.
@@
We define the proof and disproof numbers so that they give lower bounds on the size of the proof of a given position, which guarantees that PNS finds a minimal proof. For leaves, we initialize them to $1$. For atomic nodes, the proof and disproof numbers and the next node $w$ are defined as in the original setting. If $u$ is a Grundy node, then it corresponds to an atomic position $Q$ and we use $v_{i}$ to denote the child of $u$ representing $Q+\ast i$. The children of $u$ are atomic and are generated incrementally with increasing $i$, each only after the previous one has been proved, as prescribed by the recursive procedure for the calculation of $gn(Q)$. If $v_{j}$ is the last child of $u$ generated so far, then we set $pn(u)=dn(u)=\min{\{pn(v_{j}),dn(v_{j})\}}$ and $w=v_{j}$. Finally, let $v$ be a decomposable node. Then $v$ corresponds to a couple $P+\ast n$ where $P=P_{1}+\cdots+P_{k}$ and $k\geq 2$. We use $u_{1},\dots,u_{k}$ to denote the nodes where each $u_{i}$ is a Grundy node corresponding to $P_{i}$ and $v^{\prime}_{k}$ is an atomic node corresponding to $P_{k}+\ast n^{\prime}$. Note that only $u_{1},\dots,u_{k-1},v^{\prime}_{k}$ are the children of $v$, and that in order to generate $v^{\prime}_{k}$, we first need to know the Grundy numbers of $P_{1},\dots,P_{k-1}$, as $n^{\prime}=n\oplus gn(P_{1})\oplus\cdots\oplus gn(P_{k-1})$. Thus, if $v^{\prime}_{k}$ has not been generated, we use $u_{k}$ instead and set $pn(v)=dn(v)=\sum_{i=1}^{k}pn(u_{i})$ and $w$ to be the first node from $u_{1},\dots,u_{k-1}$ with unknown Grundy number. Otherwise, we let $pn(v)=pn(v^{\prime}_{k})$, $dn(v)=dn(v^{\prime}_{k})$, and $w=v^{\prime}_{k}$.
We improve this version of PNS from Lemoine (2011) by following Schiff et al. (1994) and merging nodes representing identical states to transform the tree into a directed acyclic graph. This significantly reduces the computation time without increasing memory overhead when combined with storing all touched nodes.
@@
To simplify the definition of the thresholds, we replace each Grundy node $u$ in its decomposable parent with the node $v_{j}$, which would be selected as the next node if we were in $u$. Then, we can keep the thresholds only for atomic and decomposable nodes. Now, let $v$ be the currently visited node, and let $w$ be the next node to be selected. For an atomic $v$, we let $w^{\prime}$ be its child with the second-lowest disproof number, and we set the thresholds as
@@
where ps(w)=ds(v)+dn(v)pn(w)ps(w)=ds(v)+dn(v)-pn(w) and $ds(w)=ps(v)$. For a decomposable $v$, we set $pt(w)=pt(v)$, $dt(w)=dt(v)$, and $mt(w)=mt(v)$ if $v^{\prime}_{k}$ has been generated. If $v^{\prime}_{k}$ has not been generated yet, we define the thresholds as
@@
where $ps(w)=ds(w)=0$ and $t(v)=\min\{pt(v),dt(v),\allowbreak mt(v)-\min\{ps(v),ds(v)\}\}$. For the root $r$, we initialize $pt(r)=dt(r)=mt(r)=\infty$ and $ps(r)=ds(r)=0$. With this choice of thresholds, we obtain the following result.
@@
See Appendix for the proof. To balance the time complexity and memory consumption of DFPN with GN, we also incorporate a transposition table of proof and disproof numbers, maintained by the replacement strategy proposed by Nagai (2002). In addition, we store the Grundy numbers derived during computation in a separate database. Since this database is much smaller and fits easily into memory, we do not apply any replacement strategy to it.
@@
Our new parallel variant PNS-PDFPN of Proof-Number Search operates on two parallelized levels. The first PNS level targets distributed-memory systems and is based on Job-Level Proof-Number Search used by Schaeffer et al. (2007), Saffidine et al. (2012), and Wu et al. (2013). At this level, the master process maintains the current proof state and repeatedly assigns jobs to workers for asynchronous second-level processing. At the second level, each worker performs the parallel DFPN algorithm by Kaneko (2010), which we call PDFPN, to utilize shared memory within a single cluster node. We also introduce the sharing of key results between workers to reduce search overhead. We now describe both levels in detail.
@@
The master and workers typically run on separate nodes of a cluster and communicate over an interconnected network. When a worker becomes idle, the master selects a so-called pseudo-MPN leaf $\ell$ and assigns it to the worker. The worker then processes $\ell$ asynchronously until it is solved or the maximum number of expansions is reached. The resulting proof and disproof numbers of $\ell$ and its children are then sent back to the master, which uses them to expand $\ell$ in the master tree and updates its proof and disproof numbers. During processing of $\ell$, the worker periodically sends updated values $pn(\ell)$ and $dn(\ell)$ to keep the master tree as up-to-date as possible. To prevent reassignment of the same jobs, the assigned leaves are locked, and the definition of proof and disproof numbers is slightly adjusted to avoid misleading attraction to the locked leaves; see Wu et al. (2013) for details.
Similar to Liang et al. (2015) and their JL-UCT Search, our master maintains a database of key computed results shared with all workers to reduce search overhead. Newly derived results are sent to workers along with their assigned jobs, while workers report new results back to the master with each update. Finally, if multiple workers are located on the same node, then they are grouped together to share the same local version of the database.
@@
Our workers are based on DFPN to make full use of available memory; thus, we parallelize them using PDFPN. In this algorithm, there are as many threads as the number of cores assigned to the worker. Each thread performs an independent DFPN search in the subtree of $\ell$, while sharing the lock-protected transposition table of proof and disproof numbers. To decrease redundant thread computation in shared subtrees, the disproof numbers are virtually increased during the selection of the next node $w$ by the number of threads currently computing in the corresponding subtrees. The thresholds in (1) and (2) are then modified by subtracting $th(w)$ from the term $dn(w^{\prime})+1$, where $th(w)$ is the number of threads in the subtree of $w$. Additionally, if a thread solves a position currently computed by a different one, the other thread is notified to backtrack to that position.